# Micellization in active matter of asymmetric self-propelled particles: Experiments

### v2

@@ VI Discussion
In this work, we experimentally demonstrate the emergence of micellization in self-propelled particle ensembles implemented by robotic swarms, which is governed by an interplay of particles’ activity and their shape asymmetry, in contrast to micellization of surfactants in water solutions due to their inability to form hydrogen bonds Némethy and Scheraga (1962a, b, c). To prevent the dominance of boundary condensation Giomi et al. (2013); Deblais et al. (2018) which would inevitably mask bulk phenomena such as micelle formation, we consider two scenarios: systems of chiral robots moving on a flat surface along circular orbits with diameters much larger than the size of an individual robot, yet considerably smaller than the diameter of the boundary, and non-chiral robots moving linearly but placed in a parabolic satellite dish creating a soft localizing potential.
As we demonstrate, the presence or absence of micellization is defined by the location of the robot center of inertia, which supports the numerical predictions Kruglov and Borisov (2021). In particular, the robots can form micelles if their center of inertia is shifted towards the nose, while the micellization is suppressed for the center of inertia located close to the geometrical center. Such a behavior is clearly observed in both cases, for swarms of chiral robots on a flat surface and for swarms of non-chiral robots in a parabolic potential. However, the presence of a localizing potential reduces the characteristic time of micelle formation and greatly enhances the mean lifetime of micelles. The stability of the micelles also increases with increasing friction between individual robots.
To underpin the microscopic physical mechanisms responsible for the formation of micelles upon changing the location of robot center of inertia, we performed two series of experiments addressing the motion characteristics of individual robots and the properties of two-robot collisions. Although the observed difference in the single-robot behavior cannot be directly linked to the emergence of micellization, our two-robot collision experiments clearly indicate the formation of two different types of stable two-robot clusters depending on the location of robot center of inertia. One of them represents a quarter of the micelle, thus being its constructive block. Therefore, the observed emergence of micellization is at least a two-particle effect.
@@
The second direction for further research is to consider swarms with non-zero net chirality in contrast to those studied in Sec. IV.2. For chiral particles, the emergence of edge states has been reported in physical settings involving swarms of various constituents ranging from self-rotating Yang et al. (2020) or self-propelled Barois et al. (2020) robots to bacteria with small net chirality Li et al. (2024) or even dancing pairs of people Du et al. (2024). In the context of robotic swarm micellization, the presence of such edge states could modify the rate of robot condensation at the boundary. Moreover, fully chiral micelles with a considerable angular momentum are expected to demonstrate stability properties and formation mechanisms different from those of non-chiral or slightly chiral micelles.
Another prospective direction is the implementation of the reported micellization at smaller scales, for example, with millimeter-size particles propelled by Marangoni flows Shabaniverki et al. (2021); Kumar et al. (2022), and at the microscale, where new interaction mechanisms come into play, such as increased Van-der-Waals forces. In terms of microscale implementations, Janus particles appear to be the most promising, as they allow one to change the activity of particles by an external illumination Villa and Pumera (2019); Palacci et al. (2013); Dai et al. (2016). Moreover, self-propelled Janus particles with various asymmetric shapes have been synthesized Dai et al. (2016); Shah et al. (2021); Liu et al. (2022), making the proposed design feasible.
@@ Acknowledgements.
We acknowledge fruitful discussions with Dr. Timofey Kruglov, who brought the idea of active matter micellization to our attention, Dr. Alexander Borisov, and Prof. Anton Souslov. The work is supported by the Russian Science Foundation (project 24-79-10314).References
@@
- Li et al. (2021) Shengkai Li, Bahnsikha Dutta, Sarah Cannon, Joshua Daymude, Ram Avinery, Enes Aydin, Andréa Richa, Daniel Goldman, and Dana Randall, “Programming active cohesive granular matter with mechanically induced phase changes,” Science Advances 7 (2021), 10.1126/sciadv.abe8494.
@@
- Liu et al. (2022) Jingjing Liu, Zheng-wei Li, Ze-Guo Chen, Weiyuan Tang, An Chen, Bin Liang, Guancong Ma, and Jian-Chun Cheng, “Experimental realization of weyl exceptional rings in a synthetic three-dimensional non-hermitian phononic crystal,” Physical Review Letters 129 (2022), 10.1103/PhysRevLett.129.084301.
- Dmitriev et al. (2023) Alexey A. Dmitriev, Alina D. Rozenblit, Vadim A. Porvatov, Mikhail K. Buzakov, Anastasia A. Molodtsova, Daria V. Sennikova, Vyacheslav A. Smirnov, Oleg I. Burmistrov, Timur I. Karimov, Ekaterina M. Puhtina, and Nikita A. Olekhno, “Swarmodroid 1.0: A modular bristle-bot platform for robotic active matter studies,” (2023), arXiv:2305.13510 [cond-mat.soft] .
@@
- (38) See Supplemental Material at [URL] for (i) analytical analysis of the properties of circulangles; (ii) derivation of the order parameter; (iii) the details of numerical simulations; (iv) description of the experimental setup; and (v) discussion of the experiments with large friction between robots .

### v3

@@ VI Discussion
In this work, we experimentally demonstrate the emergence of micellization in self-propelled particle ensembles implemented by robotic swarms, which is governed by an interplay of particles’ activity and their shape asymmetry, in contrast to micellization of surfactants in water solutions due to their inability to form hydrogen bonds Némethy and Scheraga (1962a, b, c). To prevent the dominance of boundary condensation Giomi et al. (2013); Deblais et al. (2018) which would inevitably mask bulk phenomena such as micelle formation, we consider two scenarios: systems of chiral robots moving on a flat surface along circular orbits with diameters much larger than the size of an individual robot, yet considerably smaller than the diameter of the boundary, and nonchiral robots moving linearly but placed in a parabolic satellite dish creating a soft localizing potential.
As we demonstrate, the presence or absence of micellization is defined by the location of the robot center of inertia, which supports the numerical predictions Kruglov and Borisov (2021). In particular, the robots can form micelles if their center of inertia is shifted towards the nose, while the micellization is suppressed for the center of inertia located close to the geometrical center. Such a behavior is clearly observed in both cases, for swarms of chiral robots on a flat surface and for swarms of nonchiral robots in a parabolic potential. The presence of such a localizing potential reduces the characteristic time of micelle formation and greatly enhances the mean lifetime of micelles. However, such micellelike clusters in the considered system can break apart due to the interaction with single robots or other clusters, and are generally metastable, in contrast to micelles formed by surfactants. As a result, a transition to the fully micellar phase in the considered system appears unrealistic. Additional studies of the stability of individual micelles as well as their interaction with the rest of the robotic swarm are discussed in the supplementary material Sup .
To underpin the microscopic physical mechanisms responsible for the formation of micelles on changing the location of the robot center of inertia, we performed two series of experiments addressing the motion characteristics of individual robots and the properties of two-robot collisions. Although the observed difference in the single-robot behavior cannot be directly linked to the emergence of micellization, our two-robot collision experiments clearly indicate the formation of two different types of stable two-robot clusters depending on the location of robot center of inertia. One of them represents a quarter of the micelle, thus being its constructive block. Therefore, the observed emergence of micellization is at least a two-particle effect.
@@
The second direction for further research is to consider swarms with nonzero net chirality in contrast to those studied in Sec. IV.2. For chiral particles, the emergence of edge states has been reported in physical settings involving swarms of various constituents ranging from self-rotating Yang et al. (2020) or self-propelled Barois et al. (2020) robots to bacteria with small net chirality Li et al. (2024) or even dancing pairs of people Du et al. (2024). In the context of robotic swarm micellization, the presence of such edge states could modify the rate of robot condensation at the boundary. Moreover, fully chiral micelles with a considerable angular momentum are expected to demonstrate stability properties and formation mechanisms different from those of nonchiral or slightly chiral micelles.
Development of an analytical theoretical description for the reported micellization is yet another direction for further research. The system considered in the present paper is inherently nonequilibrium as the robots constantly convert the energy stored in their batteries to kinetic energy. It is also not clear whether an effective temperature can be properly defined in such a system Loi et al. (2008); Takatori and Brady (2015); C.P. and Joy (2022); Boudet et al. (2022); Caprini et al. (2024). This renders the approach of considering micellization as a reversible chemical reaction at equilibrium Phillips (1955); Nagarajan (2014) (which is conventionally used to derive the free energy of micellization) inapplicable to our system. Moreover, in surfactant solutions, the micelle concentration increases as more surfactant is added to the mixture, in accordance with Le Chatelier’s principle. However, for the reported micellization, increasing the robot concentration in some cases instead leads to a decrease in the number of micellelike clusters observed for abrasive -robots, as is discussed in detail in the supplementary material Sup . Moreover, a further increase in the number of robots is expected to completely prevent micellization due to the jamming transition Barois et al. (2019).
Another prospective direction is the implementation of the reported micellization at smaller scales, for example, with millimeter-size particles propelled by Marangoni flows Shabaniverki et al. (2021); Kumar et al. (2022), and at the microscale, where new interaction mechanisms come into play, such as increased van der Waals forces. In terms of microscale implementations, Janus particles appear to be the most promising, as they allow one to change the activity of particles by an external illumination Villa and Pumera (2019); Palacci et al. (2013); Dai et al. (2016). Moreover, self-propelled Janus particles with various asymmetric shapes have been synthesized Dai et al. (2016); Shah et al. (2021); Liu et al. (2022), making the proposed design feasible.
@@ Acknowledgements.
We acknowledge fruitful discussions with Dr. Timofey Kruglov, who brought the idea of active - matter micellization to our attention; Dr. Alexander Borisov, Prof. Mikhail Lapine; and Prof. Anton Souslov. The work is supported by the Russian Science Foundation (project 24-79-10314).References
@@
- Li et al. (2021) Shengkai Li, Bahnsikha Dutta, Sarah Cannon, Joshua Daymude, Ram Avinery, Enes Aydin, André’a Richa, Daniel Goldman, and Dana Randall, “Programming active cohesive granular matter with mechanically induced phase changes,” Science Advances 7, eabe8494 (2021).
@@
- Liu et al. (2022) Meihuan Liu, Jiamiao Jiang, Haixin Tan, Bin Chen, Juanfeng Ou, Hong Wang, Jia Sun, Lu Liu, Fei Wang, Junbin Gao, Chang Liu, Fei Peng, Yun Liu, and Yingfeng Tu, “Light-driven Au–ZnO nanorod motors for enhanced photocatalytic degradation of tetracycline,” Nanoscale 14, 12804–12813 (2022).
- Dmitriev et al. (2023) Alexey A. Dmitriev, Alina D. Rozenblit, Vadim A. Porvatov, Mikhail K. Buzakov, Anastasia A. Molodtsova, Daria V. Sennikova, Vyacheslav A. Smirnov, Oleg I. Burmistrov, Timur I. Karimov, Ekaterina M. Puhtina, and Nikita A. Olekhno, “Swarmodroid 1.0: A modular bristle-bot platform for robotic active matter studies,” (2023), updated versions of the AMPy code and its documentation will be made available online (https://github.com/swarmtronics/ampy), arXiv:2305.13510 .
@@
- (31) See Supplemental Material at http://link.aps.org/supplemental/10.1103/4yzc-ppgy for (i) a description of the experimental setup; (ii) details on the choice of an abrasive material for robot side surfaces; (iii) detailed studies of a single-robot motion in a parabolic potential; (iv) experiments on a single-robot motion under a rotational excitation; (v) studies of two-robot collisions with different parameters; (vi) the classification of various several-robot clusters; (vii) studies of the effects of aggregation induced by a parabolic potential; and (viii) a comparison with micellization of surfactants. Additionally, Supplemental Video 1 demonstrates the condensation of robots at a vertical wall; Supplemental Video 2 demonstrates micellization at a flat surface with a vertical wall; Supplemental Video 3 demonstrates micellization in a parabolic potential; Supplemental Video 4 demonstrates the motion of a single robot in the parabolic potential; Supplemental Video 5 demonstrates the formation of stable two-robot clusters; Supplemental Video 6 demonstrates an example of a stabilized micelle; Supplemental Video 7 illustrates the stability of an isolated micelle like cluster; and Supplemental Video 8 shows the stability of artificially assembled micelle like cluster .
- Scholz et al. (2016) Christian Scholz, Sean D’Silva, and Thorsten Pöschel, “Ratcheting and tumbling motion of vibrots,” New Journal of Physics 18, 123001 (2016).
- Porvatov et al. (2021) Vadim A. Porvatov, Alina D. Rozenblit, Alexey A. Dmitriev, Oleg I. Burmistrov, Daria A. Petrova, Georgy Yu. Gritsenko, Ekaterina M. Puhtina, Egor I. Kretov, Dmitry S. Filonov, Anton Souslov, and Nikita A. Olekhno, “Optimizing self-rotating bristle-bots for active matter implementation with robotic swarms,” Journal of Physics: Conference Series 2086, 012202 (2021).
@@
- Loi et al. (2008) Davide Loi, Stefano Mossa, and Leticia F Cugliandolo, “Effective temperature of active matter,” Physical Review E 77, 051111 (2008).
- Takatori and Brady (2015) Sho C. Takatori and John F. Brady, “Towards a thermodynamics of active matter,” Physical Review E 91, 032117 (2015).
- C.P. and Joy (2022) Sanjay C.P. and Ashwin Joy, “Effective temperature and einstein relation for particles in active matter flows,” Physical Review E 105, 065114 (2022).
- Caprini et al. (2024) Lorenzo Caprini, Davide Breoni, Anton Ldov, Christian Scholz, and Hartmut Löwen, “Dynamical clustering and wetting phenomena in inertial active matter,” Communications Physics 7, 343 (2024).
- Phillips (1955) J.N. Phillips, “The energetics of micelle formation,” Transactions of the Faraday Society 51, 561–569 (1955).
- Nagarajan (2014) Ramanathan Nagarajan, “One hundred years of micelles: evolution of the theory of micellization,” in Surfactant Science and Technology: Retrospects and Prospects, edited by Laurence S. Romsted (Taylor and Francis Boca Raton, FL, 2014) pp. 3–52.